# Sign-In to the Lottery: Reparameterizing Sparse Training From Scratch

### v1

@@ 7 Conclusion
First we show that for balanced initializations $|a_{in}|=w_{in}>0$ enter the region $\Gamma_{0}$. In case $a_{in}=w_{in}>0$, we have $(a_{in},w_{in})\in\Gamma_{0}$. In case $-a_{in}=w_{in}>0$, we have that $(a_{in},w_{in})\in\bar{\Gamma}_{0}\setminus\Gamma_{0}$ i.e. the boundary of $\Gamma_{0}$. Therefore, we need to show the gradient field at $(a_{in},w_{in})$ points into $\Gamma_{0}$.
@@
Appendix B Algorithm Initialization
@@
Appendix C Multiple Neurons Visualization
@@
Appendix D Training Details
In this section we provide the details of the additional hyperparameters used for experiments. Table 6 we provide the standard hyperparameters such as learning rate and weight decay. We note that the loss used for both CIFAR10 and CIFAR100 is the cross-entropy loss and for ImageNet the label smoothing loss with confidence $0.9$. For the $m\odot w$ parametrization we also apply Frobenius decay i.e. $(m\odot w)^{2}$ with the same strength as weight decay for CIFAR10 and CIFAR100. We provide these extra details in Table 7.
@@ Appendix D Training Details
Dataset Model LR Weight Decay Epochs Batch Size Optim LR Schedule
CIFAR10 ResNet20 $0.2$ $41e-$ 4 $150$ $512$ SGD, $m=0.9$ Triangular
CIFAR100 ResNet18 $0.2$ $41e-$ 4 $150$ $512$ SGD, $m=0.9$ Triangular
ImageNet ResNet50 $0.25$ $55e-$ 5 $100$ $1024$ SGD, $m=0.9$ Triangular
DeIT Small $0.001$ $11e-$ 1 $300$ $1024$ AdamW Triangular
@@
CIFAR10 ResNet20 $1$ $75$ $41e-$ 4 00
CIFAR100 ResNet18 $1$ $75$ $41e-$ 4 00
ImageNet ResNet50 $1$ $75$ 00 $55e-$ 5
DeIT Small $1$ $225$ 00 $11e-$ 1
Appendix E Experiments on two-layer networks with Sign-In
We empirically verify the ability to recover correct signs for dense-to-sparse methods and Sign-In on the two-layer network with a student-teacher setup described in Section 5. Further, we also empirically evaluate the effect of combining dense-to-sparse methods with Sign-In on the two-layer network. We repeat $100$ runs for each case, with data generated from a teacher $\tilde{a}=1,\tilde{w}=\left[1,0,..,0\right]$, such that the labels are $y=\tilde{f_{0}}(\tilde{a},\tilde{w},z)$. For the multidimensional case we use $d=5$. Each setting is trained with Gradient Descent for $52e$ 5 iterations with a learning rate $=0.01$. Results reported in Table 8 are the success fraction over $100$ runs. We empirically observe that it is possible to successfully recover all cases when combining overparameterized training with Sign-In. Even the most difficult case that is unreachable for both on their own i.e. $a<0$ and $w<0$. Now it is possible due to Sign-In being able to flip $a$, whereas overparameterized training can flip $w$ after that. This is also observed with improved performance of AC/DC with Sign-In (see Table 5).
Table 8: Experiments on two-layer networks in the student-teacher setup presented in Theorem 5.2.
@@ Appendix E Experiments on two-layer networks with Sign-In
Sparse $1$ 00 00 00
Overparam $1$ 00 $1$ 00
Sign-In $1$ $1$ 00 00
Overparam + Sign-In $1$ $1$ $1$ $1$
@@
Appendix F Sign-In flips signs in later layers
Here we validate that the sign flips enabled by Sign-In are different from those enabled with dense-to-sparse training. For the case $s=0.9$ reported in Table 5 for AC/DC and AC/DC + Sign-In, we plot the layerwise sign flips for three stages, initialization to warmup ($10$-th epoch), warmup to final and initialization to final. Additionally, we also compare the layerwise sign flips of sparse training a randomly reinitialized AC/DC mask. While AC/DC identifies the correct sign flips upto warmup, Sign-In enables more sign flips during subsequent training for later layers in the network (see orange line in Figure 8(b)).
@@ Appendix F Sign-In flips signs in later layers
Appendix G Sign-In with other PaI methods
We report sparse training performance with Sign-In across other PaI methods including SNIP (Lee et al., 2019) and Synflow (Tanaka et al., 2020) in Table 9. We use the same setup as described in Appendix D.
@@ Appendix G Sign-In with other PaI methods
First, we observe that in case of masks identified by STR and PaI with Snip and Synflow achieve poor performance for both standard training and Sign-In. The layerwise sparsity ratios of these masks reveal sparse initial layers ($>90\%$) or highly sparse ($>99\%$) later layers, which potentially affect the overall trainability of the network. In contrast, the RiGL mask, which is a random mask with balanced layerwise sparsity ratio (Gadhikar et al., 2023), performs the best over all other masks with Sign-In. Here initial layers are dense while later layers always have sparsity $<99\%$.
Second, we observe that when initial layers have low sparsity, Sign-In can outperform standard training (RiGL, AC/DC, Snip) but struggles if this is not the case (STR, Synflow). A potential cause would be the inability of initial layers to flip $w$ (see Theorem 5.2).
@@
Appendix H Rescaling $m\odot w$ in Sign-In
While our proposed pointwise reparameterization with $m\odot w$ allows better sign flips for sparse training, we find that the rescaling step outlined in Algorithm 1 is essential to maximize the benefits of the reparameterization as show with improved performance of Sign-In over $m\odot w$ in Table 11.
@@ Appendix H Rescaling $m\odot w$ in Sign-In
Appendix I Code for Sign-In in PyTorch
@@
import torch.nn as nn
class ConvMaskMW(nn.Conv2d):
@@
**kwargs: Keyword arguments for nn.Conv2d.
@@
self.m = nn.Parameter((u + alpha/u)/np.sqrt(2), requires_grad=True)

### v2

@@ 8 Acknowledgements
First we show that for balanced initializations $|a_{in}|=w_{in}>0$ enter the region $\Gamma_{0}$. In case $a_{in}=w_{in}>0$, we have $(a_{in},w_{in})\in\Gamma_{0}$. In case $a_{in}=w_{in}>0$, we have that $(a_{in},w_{in})\in\bar{\Gamma}_{0}\setminus\Gamma_{0}$ i.e. the boundary of $\Gamma_{0}$. Therefore, we need to show the gradient field at $(a_{in},w_{in})$ points into $\Gamma_{0}$.
@@
Appendix B Sign-In algorithm
We provide pseudocode for Sign-In for in Section˜4 with rescaling as explained in Section˜4.
Algorithm 1 Sign-In
for $i=1$ to $T$ do
  if $i$ mod $p==0$ and $i<T_{2}$ then
  end if
end for
Return Model parameters $\theta_{T}=S\odot m_{T}\odot w_{T}$
Appendix C Algorithm Initialization
@@
Appendix D Multiple Neurons Visualization
@@
Appendix E Training Details
In this section we provide the details of the additional hyperparameters used for experiments. Table 4 we provide the standard hyperparameters such as learning rate and weight decay. We note that the loss used for both CIFAR10 and CIFAR100 is the cross-entropy loss and for ImageNet the label smoothing loss with confidence $0.9$. For the $m\odot w$ parametrization we also apply Frobenius decay i.e. $(m\odot w)^{2}$ with the same strength as weight decay for CIFAR10 and CIFAR100. We provide these extra details in Table 5. The code used is based on TurboPrune as in (Nelaturu et al., 2024).
@@ Appendix E Training Details
Dataset Model LR WD Epochs Batch Size Optim Schedule
CIFAR10 ResNet20 $0.2$ 1e41e-4 $150$ $512$ SGD Triangular
CIFAR100 ResNet18 $0.2$ 1e41e-4 $150$ $512$ SGD Triangular
ImageNet ResNet50 $0.25$ 5e55e-5 $100$ $1024$ SGD Triangular
DeIT Small $0.001$ 1e11e-1 $300$ $1024$ AdamW Triangular
@@
CIFAR10 ResNet20 11 $75$ 1e41e-4 0
CIFAR100 ResNet18 11 $75$ 1e41e-4 0
ImageNet ResNet50 11 $75$ 0 5e55e-5
DeIT Small 11 $225$ 0 1e11e-1
Appendix F Experiments on two-layer networks with Sign-In
We empirically verify the ability to recover correct signs for dense-to-sparse methods and Sign-In on the two-layer network with a student-teacher setup described in Section˜5. Further, we also empirically evaluate the effect of combining dense-to-sparse methods with Sign-In on the two-layer network. We repeat $100$ runs for each case, with data generated from a teacher $\tilde{a}=1,\tilde{w}=\left[1,0,..,0\right]$, such that the labels are $y=\tilde{f_{0}}(\tilde{a},\tilde{w},z)$. For the multidimensional case we use $d=5$. Each setting is trained with Gradient Descent for 2e52e-5 iterations with a learning rate =0.01=0.01. Results reported in Table˜6 are the success fraction over $100$ runs. We empirically observe that it is possible to successfully recover all cases when combining overparameterized training with Sign-In. Even the most difficult case that is unreachable for both on their own i.e. $a<0$ and $w<0$. Now it is possible due to Sign-In being able to flip $a$, whereas overparameterized training can flip $w$ after that. This is also observed with improved performance of AC/DC with Sign-In (see Table˜11).
Table 6: Experiments on two-layer networks in the student-teacher setup presented in Theorem˜5.2.
@@ Appendix F Experiments on two-layer networks with Sign-In
Sparse 11 0 0 0
Overparam 11 0 11 0
Sign-In 11 11 0 0
Overparam + Sign-In 11 11 11 11
@@
Appendix G Sign-In flips signs in later layers
Here we validate that the sign flips enabled by Sign-In are different from those enabled with dense-to-sparse training. For the case $s=0.9$ reported in Table˜11 for AC/DC and AC/DC + Sign-In, we plot the layerwise sign flips for three stages, initialization to warmup ($10$-th epoch), warmup to final and initialization to final. Additionally, we also compare the layerwise sign flips of sparse training a randomly reinitialized AC/DC mask. While AC/DC identifies the correct sign flips upto warmup, Sign-In enables more sign flips during subsequent training for later layers in the network (see orange line in Figure˜7(b)).
@@ Appendix G Sign-In flips signs in later layers
Appendix H Sign-In with other PaI methods
We report sparse training performance with Sign-In across other PaI methods including SNIP (Lee et al., 2019) and Synflow (Tanaka et al., 2020) in Table˜7. We use the same setup as described in Appendix E.
@@ Appendix H Sign-In with other PaI methods
First, we observe that in case of masks identified by STR and PaI with Snip and Synflow achieve poor performance for both standard training and Sign-In. The layerwise sparsity ratios of these masks reveal sparse initial layers ($90\%$) or highly sparse ($99\%$) later layers, which potentially affect the overall trainability of the network. In contrast, the RiGL mask, which is a random mask with balanced layerwise sparsity ratio (Gadhikar et al., 2023), performs the best over all other masks with Sign-In. Here initial layers are dense while later layers always have sparsity $99\%$.
Second, we observe that when initial layers have low sparsity, Sign-In can outperform standard training (RiGL, AC/DC, Snip) but struggles if this is not the case (STR, Synflow). A potential cause would be the inability of initial layers to flip $w$ (see Theorem˜5.2).
@@
Appendix I Sign Importance
In order to identify the amount of correct signs required for training with the sparse mask to match the respective dense-to-sparse baseline (via sign-alignment), we perturb an increasing fraction of signs as reported in Table 9. Perturbing more than $10\%$ of the signs starts degrading performance rapidly.
Table 9: Ablation study on the importance of weight signs.
Method $s=0.9$
AC/DC baseline (from Table 1) $74.68$
AC/DC Mask + $100\%$ correct signs (from Table 1) $74.88$
AC/DC Mask + $95\%$ correct signs $74.44$
AC/DC Mask + $90\%$ correct signs $73.54$
AC/DC Mask + $80\%$ correct signs $71.81$
AC/DC Mask + random signs (from Table 1) $70.88$
Appendix J Vision transformer with Sign-In
Sign-In is also able to improve sparse training of a randomly masked vision transformer. A DeiT Small architecture (trained for $300$ epochs from scratch) gains more than $1\%$ in performance with Sign-In, as reported in Table˜10.
Table 10: DeiT Small on ImageNet, trained from scratch.
Method $s=0.5$
Random $69.31(\pm 0.7)$
Random + Sign-In (ours) $\mathbf{70.62(\pm 0.4)}$
Appendix K Sign-In improves other sparse training methods
Only a combination of Sign-In and overparameterization is able to solve all sign flip challenges in our theoretical setting, as also confirmed by numerical experiments, which are provided in Table˜6 in Appendix˜F. Therefore, we expect the combination of dense-to-sparse methods with Sign-In to achieve further gains in performance. Table˜11 confirms that AC/DC improves with Sign-In. However, the improvement is less significant compared to PaI.
Table 11: Sign-In improves AC/DC for a ResNet50 trained on ImageNet.
Method $s=0.8$ $s=0.9$ $s=0.95$
AC/DC $75.83(\pm 0.02)$ $74.75(\pm 0.02)$ $72.59(\pm 0.11)$
+ Sign-In
Sign-In also improves dynamic sparse training with RiGL and MEST by enabling better sign flips.
Table 12: Sign-In improves RiGL for a ResNet50 trained on ImageNet.
Method $s=0.8$ $s=0.9$ $s=0.95$
RiGL $75.02(\pm 0.1)$ $73.7(\pm 0.2)$ $71.89(\pm 0.07)$
RiGL $75.02(\pm 0.1)$ $\mathbf{74.27(\pm 0.08)}$ $\mathbf{73.07(\pm 0.17)}$
+ Sign-In
Table 13: Sign-In improves MEST for a ResNet50 trained on ImageNet.
Method $s=0.8$ $s=0.9$ $s=0.95$
MEST $74.67$ $73.1$ $70.84$
MEST $\mathbf{74.74}$ $\mathbf{73.32}$ $\mathbf{71.41}$
+ Sign-In
Appendix L Rescaling in Sign-In
While our proposed pointwise reparameterization with $m\odot w$ allows better sign flips for sparse training, we find that the rescaling step outlined in Appendix˜B is essential to maximize the benefits of the reparameterization as show with improved performance of Sign-In over $m\odot w$ in Table˜14.
@@ Appendix I Sign Importance
Appendix M Code for Sign-In in PyTorch
@@
import torch.nn as $n$
class ConvMaskMW($n$.Conv2d):
@@
**kwargs: Keyword arguments for $n$.Conv2d.
@@
self.m = $n$.Parameter((u + alpha/u)/np.sqrt(2), requires_grad=True)
@@
Appendix N FLOPS required for Sign-In
We perform a FLOP comparison for Sign-In vs. other PaI methods. The FLOP overhead due to the reparameterization in Sign-In is negligible and only required during training, as $m\odot w$ can be merged after training, during inference.
For example, the FLOPs for a convolutional layer are:
$(2\times H_{\text{out}}\times W_{\text{out}}\times C_{\text{out}}\times K\times K\times C_{\text{in}})+(C_{\text{out}}\times C_{\text{in}}\times K\times K)$
An additional term $(C_{\text{out}}\times C_{\text{in}}\times K\times K)$ is introduced by the elementwise product $m\odot w$, but it is negligible with respect to the other term.
We report the FLOPs per forward pass for a ResNet50.
Table 15: FLOP Comparison of Sign-In vs. other PaI methods on ResNet50. Sign-In introduces minimal overhead while improving accuracy.
Method Training FLOPs (G) Inference FLOPs (G) Acc
Dense 8.2 8.2 76.89
Sparse ($s=0.8$) 4.99 4.99 73.87
Sign-In ($s=0.8$) 5 4.99 74.12
Sparse ($s=0.9$) 3.89 3.89 71.56
Sign-In ($s=0.9$) 3.9 3.89 72.19
Appendix O High sparsity regime ablation
In this section we provide results for the high sparsity regime.
Table 16: Sign-In with PaI methods on CIFAR10 and CIFAR100 in the high sparsity regime.
Dataset Method $s=0.98$ $s=0.99$
CIFAR10 + ResNet20 Random $88.41(\pm 0.23)$ $86.41(\pm 0.1)$
Random + Sign-In $89.37(\pm 0.14)$ $\mathbf{87.83(\pm 0.11)}$
SNIP $88.12(\pm 0.5)$ $85.82(\pm 0.2)$
SNIP + Sign-In $\mathbf{89.88(\pm 0.09)}$ $87.27(\pm 0.13)$
Synflow $88.15(\pm 0.17)$ $85.32(\pm 0.24)$
Synflow + Sign-In $89.26(\pm 0.05)$ $86.47(\pm 0.06)$
CIFAR100 + ResNet18 Random $74.01(\pm{0.1})$ $73.07(\pm{0.19})$
Random + Sign-In $\mathbf{75.32(\pm 0.25)}$ $\mathbf{73.94(\pm 0.13)}$
SNIP $74.08(\pm 0.28)$ $72.73(\pm 0.32)$
SNIP + Sign-In $75.16(\pm 0.15)$ $73.64(\pm 0.32)$
Synflow $73.41(\pm 0.28)$ $71.37(\pm 0.42)$
Synflow + Sign-In $75.16(\pm 0.20)$ $73.64(\pm 0.33)$